Suppose a bank’s liquidity division estimates that it holds $19 million in hot money deposits against which it will hold an 80 percent liquidity reserve, $54 million in vulnerable funds against which it plans to hold a 25 percent reserve and $112 million in stable or core funds against which it will hold a 5 percent liquidity reserve. The bank expects its loans to grow 8 percent annually; its loans currently stand at $117 million, but have recently reached $132 million. If reserve requirements on liabilities currently stand at 3 percent, what is this depository institution’s total liquidity requirement?
1
Expert's answer
2020-05-20T19:50:52-0400

"total liquidity requirement=0.8\times(19-0.03\times19)+0.25\times(54-0.03\times54)+0.05\times(112-0.03\times112)+(132+0.08\times132-117)=14.744+13.095+5.432+25.56=58.831"